# Almost equivalences between Tamarkin category and Novikov sheaves

### v2

@@ 6.2 Tamarkin-type category
Let $M$ be a manifold. We consider the category $\mathrm{Sh}^{\mathbb{G}}(M\times\mathbb{R}_{t},\mathbb{K})$ of the equivariant $\mathbb{K}$-module sheaves on $M\times\mathbb{R}_{t}$ with respect to the discrete $\mathbb{G}$-action on the left component. We denote the subcategory spanned by the object whose microsupport contained in $T^{*}M\times\mathbb{R}\times\mathbb{R}_{\leq 0}$ by $\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K})$. We set
$\mu^{\mathbb{G}}(T^{*}M):=\mathrm{Sh}^{\mathbb{G}}_{>0}(M\times\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{G}}(M\times\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K}).$ (6.2)
Then $\mu^{\mathbb{G}}(T^{*}M)$ is defined over $\mathrm{Mod}(L_{0}^{\mathbb{G}})$.
@@
$\mu^{\mathbb{G}}(T^{*}M)\cong\mathrm{Sh}(M,\mathbb{K})\otimes\mu^{\mathbb{G}}(*).$ (6.3)
By using Theorem 5.5 and Lemma 6.1, we further have
$\mu^{\mathbb{G}}(T^{*}M)\cong_{\mathrm{a}}\mathrm{Sh}(M,\mathbb{K})\otimes\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})\cong\mathrm{Sh}(M,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})).$ (6.4)
$\mathrm{Sh}(M,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))\hookrightarrow\mathrm{Sh}(M,\mathrm{Mod}(L_{0}^{\mathbb{G}})).$ (6.5)
Since $M$ is a manifold, the category $\mathrm{Sh}(M,\mathrm{Mod}(L_{0}^{\mathbb{G}}))$ gets identified with the derived category $\mathrm{Sh}(M,L_{0}^{\mathbb{G}})$ of sheaves on $M$ whose values are $L_{0}^{\mathbb{G}}$-modules.
@@
$\mu^{\mathbb{G}}(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,L_{0}^{\mathbb{G}}).$ (6.6)
@@ 6.3 Sheaf quantizations
Let $M$ be a manifold. We denote the cotangent coordinate of $\mathbb{R}_{t}$ by $\tau$. We set
$\left\{\tau>0\right\}:=\left\{(p,(t,\tau))\in T^{*}M\times T^{*}\mathbb{R}_{t}\mathrel{}\middle|\mathrel{}\tau>0\right\}.$ (6.7)
@@
An object of $\mu^{\mathbb{G}}(T^{*}M)$ is a sheaf quantization of a Lagrangian submanifold $L$ if
@@
For the construction and properties of sheaf quantizations, see the companion paper [IK]. We denote the category of $\mu^{\mathbb{G}}(T^{*}M)$ consisting of sheaf quantizations of projection-finite end-conic Lagrangians by $\mathrm{SQ}^{\mathbb{G}}(T^{*}M)$.
@@
6.4 Variant 1: Liouville manifold
We can easily generalize the construction to the case of Liouville manifolds. Let $X$ be a Liouville manifold. We denote the category of microsheaves over $X$ by $\mu sh(X)$. We set $\mu^{\mathbb{G}}(X):=\mu sh(X)\otimes\mu^{\mathbb{G}}(*)$. We then have $\mu^{\mathbb{G}}(X)\cong_{\mathrm{a}}\mu sh(X)\otimes\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}})$.
6.5 Variant 2: Energy cutoff
@@
Let $N$ be a manifold and $\mathbb{R}_{s<a}:=(-\infty,a)$ for $a>0$. Then we run the above theory to get $\mu^{\mathbb{G}}(T^{*}N\times T^{*}\mathbb{R}_{s<a})$. We consider the subcategory spanned by doubling movies $\mu^{\mathbb{G}}_{<a}(T^{*}N)$: We say an object $\mathcal{E}\in\mathrm{Sh}^{\mathbb{G}}(N\times\mathbb{R}_{s}\times\mathbb{R}_{t})$ is a doubling movie if it satisfies
@@
for some $A\subset T^{*}M\times\mathbb{R}_{t}\times\mathbb{R}_{\tau>0}$ where
$\begin{split}AA:=&\left\{(p,s,0)\in(T^{*}M\times T^{*}\mathbb{R}_{t})\times T^{*}\mathbb{R}_{s<c}\mathrel{}\middle|\mathrel{}p\in A,s\geq 0\right\}\\ &\cup\left\{(p^{\prime},t,\tau,s,\sigma)\in T^{*}M\times T^{*}\mathbb{R}_{t}\times T^{*}\mathbb{R}_{s<c}\mathrel{}\middle|\mathrel{}(p^{\prime},t-s,\tau)\in A,s\geq 0,\tau=-\sigma\right\}.\end{split}$ (6.11)
We then have
$\mu^{\mathbb{G}}_{<a}(T^{*}N)\cong\mathrm{Sh}(N,\mathbb{K})\otimes\mu^{\mathbb{G}}_{<a}(*).$ (6.12)
@@
$\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})\hookrightarrow_{\mathrm{a}}\mu^{\mathbb{G}}_{<a}(*)$ (6.13)
@@
$\mu^{\mathbb{G}}_{<a}(*)\ni 1_{\mu,a}:=\bigoplus_{c\in\mathbb{G}}\mathbb{K}_{\left\{(s,t)\mathrel{}\middle|\mathrel{}0<s<a,c\leq t<s+c\right\}}.$ (6.14)
The endmorphism of this can be easily computed as almost $L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}}$. We then have a functor
$\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})\rightarrow\mu^{\mathbb{G}}_{<a}(*);M\mapsto M\otimes 1_{\mu,a}.$ (6.15)
Then this is almost fully faithful. ∎
@@
One can see that this is orthogonal to $1_{\mu,1}$. hence $1_{\mu,1}$ does not generate the whole category. The author would like to thank T. Asano for asking a related question on the draft version.
@@
$\mu^{\mathbb{G}}_{<a}(T^{*}N)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(N,\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{a}L_{0}^{\mathbb{G}})).$ (6.17)
@@
$\mu^{\mathbb{R}}_{<a}(T^{*}N)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(N,\Lambda_{0}/T^{a}\Lambda_{0}).$ (6.18)
6.6 Variant 3: Higher-dimensional version
@@
Let $\gamma$ be a simplicial closed polyhedral proper convex cone in $\mathbb{R}^{n}$ with nonempty interior. Then $\gamma$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\gamma]$. We denote the indeterminate corresponding to $a\in\gamma$ by $T^{a}$. Let ||$\cdot$|| ⋅ | be the Euclidean norm of $\mathbb{R}^{n}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\gamma]$ generated by $T^{a}$’s with $|a|>r$ by $\mathfrak{m}(r)$. We set
@@
We say that $\gamma$ is simplicial if there exists a linear isomorphism of $\mathbb{R}^{n}$ which gives an isomorphism $\gamma\cong\mathbb{R}^{n}_{\geq 0}$.
We consider $\mathbb{R}^{n}$-equivariant sheaves on $M\times\mathbb{R}^{n}$. We denote the subcategory spanned by the object whose microsupport contained in $T^{*}M\times\mathbb{R}^{n}\times(-\gamma^{\vee})$ by $\mathrm{Sh}^{\mathbb{R}^{n}}_{\leq 0}(M\times\mathbb{R}_{t},\mathbb{K})$ where $\vee$ is polar dual. We set
$\mu^{\gamma}(T^{*}M):=\mathrm{Sh}^{\mathbb{R}^{n}}_{\mathrm{Int}(\gamma^{\vee})}(M\times\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{R}^{n}}(M\times\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{R}^{n}}_{(-\gamma^{\vee})}(M\times\mathbb{R}_{t},\mathbb{K}).$ (6.20)
By tensoring Theorem 6.3 $n$ times, we ge the following:
@@
$\mu^{\gamma}(T^{*}M)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(M,\Lambda_{0}^{\gamma}).$ (6.21)
@@ 7.1 Non-conic microsupport
Let $U$ be an open subset of $M$. Let $\phi$ be a continuous function on $U$ which is bounded below. For any connected open subset $V\subset U$, we denote the infimum value of $\phi$ by $\phi_{V}$. We define a sheaf $R^{\phi}$ as follows: For any connected open subset $V$, we set $R^{\phi}(V):=R$. For an open inclusion of connected open subsets $W\subset V$, we have $\phi_{W}\geq\phi_{V}$, we set a structure morphism $R^{\phi}(V)\rightarrow R^{\phi}(W)$ by $T^{\phi_{W}-\phi_{V}}$. Sheafifying this, we get a sheaf on $U$. We also set $R^{\phi}_{c}:=R^{\phi}\otimes_{R}R_{c}$ for $c\geq 0$.
@@
For $\mathcal{E}\in\mathrm{Mod}(R_{M})$, we set $\mu\mathrm{supp}(\mathcal{E})$ to be the closure of the complement of the following set:
$\left\{(x,\xi)\mathrel{}\middle|\mathrel{}\mathop{{\mathcal{H}}om}\nolimits(R^{\phi}_{c},\mathcal{E})_{x}\simeq 0\text{ for any $c\geq 0$ and $C^{1}$-function $\phi$ with $d\phi(x)=\xi$}\right\}.$ (7.2)
@@ 7.2 Relation to usual microsupport
Since an object of $\mathrm{Mod}(R_{M})$ is a sheaf on a manifold, we can also define the usual microsupport of [KS94].
@@
For a subset $A\subset T^{*}M$, we set
$\mathbb{R}_{\geq 0}\cdot A:=\left\{(x,\xi)\in T^{*}M\mathrel{}\middle|\mathrel{}(x,\xi^{\prime})\in A,c\in\mathbb{R}_{\geq 0}\text{ s.t.}(x,\xi)=(x,c\xi^{\prime})\right\}$ (7.4)
@@
For $\mathcal{E}\in\mathrm{Mod}(R_{M})$, we have >$SS\mathbb{R}_{>0}\cdot\mu\mathrm{supp}(\mathcal{E})\subset\SS(\mathcal{E})$.

### v3

@@ 6.2 Tamarkin-type category
Let $X$ be a manifold. We consider the category $\mathrm{Sh}^{\mathbb{G}}(X\times\mathbb{R}_{t},\mathbb{K})$ of the equivariant $\mathbb{K}$-module sheaves on $X\times\mathbb{R}_{t}$ with respect to the discrete $\mathbb{G}$-action on the right component. Again, microsupport for each object of $\mathrm{Sh}^{\mathbb{G}}(X\times\mathbb{R}_{t},\mathbb{K})$ is defined through the forgetful functor forgetting the equivariancy.
We denote the subcategory spanned by the object whose microsupport contained in $T^{*}X\times\mathbb{R}\times\mathbb{R}_{\leq 0}$ by $\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(X\times\mathbb{R}_{t},\mathbb{K})$. We set
$\mu^{\mathbb{G}}(T^{*}X):=\mathrm{Sh}^{\mathbb{G}}_{>0}(X\times\mathbb{R}_{t},\mathbb{K}):=\mathrm{Sh}^{\mathbb{G}}(X\times\mathbb{R}_{t},\mathbb{K})/\mathrm{Sh}^{\mathbb{G}}_{\leq 0}(X\times\mathbb{R}_{t},\mathbb{K}).$ (6.2)
Then $\mu^{\mathbb{G}}(T^{*}X)$ is defined over $\mathrm{Mod}(L_{0}^{\mathbb{G}})$.
@@
$\mu^{\mathbb{G}}(T^{*}X)\cong\mathrm{Sh}(X,\mathbb{K})\otimes\mu^{\mathbb{G}}(*).$ (6.3)
We then have the following:
@@
$\mu^{\mathbb{G}}(T^{*}X)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(X,L_{0}^{\mathbb{G}}).$ (6.4)
This functor will also be called $\mathfrak{A}$ in the following.
Proof.
$\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))\hookrightarrow\mathrm{Sh}(X,\mathrm{Mod}(L_{0}^{\mathbb{G}})).$ (6.5)
Since $X$ is a manifold (in particular, hypercomplete), the category $\mathrm{Sh}(X,\mathrm{Mod}(L_{0}^{\mathbb{G}}))$ gets identified with the derived category $\mathrm{Sh}(X,L_{0}^{\mathbb{G}})$ of sheaves on $X$ whose values are $L_{0}^{\mathbb{G}}$-modules. Hence, to finish the proof, it is enough to show $\mu^{\mathbb{G}}(T^{*}X)\cong_{\mathrm{a}}\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))$.
By Lemma 6.1 and Lemma 6.2 and the presentability of $\mu^{\mathbb{G}}(*)$, we have $\mu^{\mathbb{G}}(T^{*}X)\cong\mathrm{Sh}(X,\mu^{\mathbb{G}}(*))$. So, the problem is reduced to show that $\mathrm{Sh}(X,\mu^{\mathbb{G}}(*))\cong_{\mathrm{a}}\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))$.
The functor $\mathfrak{A}$ from Theorem 5.5, induces a functor $\mathrm{Sh}(X,\mu^{\mathbb{G}}(*))\rightarrow\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))$, which will also be denoted by $\mathfrak{A}$. Similarly, we consider $\mathfrak{B}\colon\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))\to\mathrm{Sh}(X,\mu^{\mathbb{G}}(*))$. For any $\mathcal{E}\in\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))$, the counit morphism $\mathcal{E}\to\mathfrak{A}\circ\mathfrak{B}(\mathcal{E})$ is an almost isomorphism. Hence $\mathfrak{A}$ is almost essentially surjective. For $\mathcal{E},\mathcal{F}\in\mathrm{Sh}(X,\mathrm{Mod}_{\mathrm{c}}(L_{0}^{\mathbb{G}}))$,
$\operatorname{Hom}(\mathfrak{A}(\mathcal{E}),\mathfrak{A}(\mathcal{F}))\cong\operatorname{Hom}(\mathfrak{B}\circ\mathfrak{A}(\mathcal{E}),\mathcal{F}).$ (6.6)
Since $\mathfrak{B}\circ\mathfrak{A}(\mathcal{E})\cong_{\mathrm{a}}\mathcal{E}$, we conclude $\operatorname{Hom}(\mathfrak{B}\circ\mathfrak{A}(\mathcal{E}),\mathcal{F})\cong_{\mathrm{a}}\operatorname{Hom}(\mathcal{E},\mathcal{F})$. This shows the almost fully faithfulness of $\mathfrak{A}$. ∎
@@ 6.3 Sheaf quantizations
Let $X$ be a manifold. We denote the cotangent coordinate of $\mathbb{R}_{t}$ by $\tau$. We set
$\left\{\tau>0\right\}:=\left\{(p,(t,\tau))\in T^{*}X\times T^{*}\mathbb{R}_{t}\mathrel{}\middle|\mathrel{}\tau>0\right\}.$ (6.7)
@@
An object of $\mu^{\mathbb{G}}(T^{*}X)$ is a sheaf quantization of a Lagrangian submanifold $L$ if
@@
For the construction and properties of sheaf quantizations, see the companion paper [IK]. We denote the full subcategory of $\mu^{\mathbb{G}}(T^{*}X)$ consisting of sheaf quantizations of projection-finite end-conic Lagrangians by $\mathrm{SQ}^{\mathbb{G}}(T^{*}X)$.
@@
6.4 Variant 1: Energy cutoff
@@
Let $X$ be a manifold and $\mathbb{R}_{s<c}:=(-\infty,c)$ for $c>0$. Then we run the above theory to get $\mu^{\mathbb{G}}(T^{*}X\times T^{*}\mathbb{R}_{s<c})$. We consider the subcategory $\mu^{\mathbb{G}}_{<c}(T^{*}X)$ consisting of the objects satisfying
@@
Obviously, we have
$\mu^{\mathbb{G}}_{<c}(T^{*}X)\cong\mathrm{Sh}(X,\mathbb{K})\otimes\mu^{\mathbb{G}}_{<c}(*).$ (6.11)
This category is a natural living place of sheaf quantizations of “obstructed” Lagrangian branes. Those are realized as doubling movie sheaf quantization of Lagrangian; objects satisfying
[…]
for some Lagrangian subset $A\subset T^{*}X\times\mathbb{R}_{t}\times\mathbb{R}_{\tau>0}$. As obstructed Lagrangian Floer complex defined over $\Lambda_{0}/T^{c}\Lambda_{0}$, the parallel story on the sheaf side can be discusses as a variant of the main story of this paper:
@@
$\mathfrak{B}_{c}\colon\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{c}L_{0}^{\mathbb{G}})\hookrightarrow_{\mathrm{a}}\mu^{\mathbb{G}}_{<c}(*)$ (6.13)
The right adjoint is given by $\mathfrak{A}_{c}:=\operatorname{Hom}(\bigoplus_{c^{\prime}\in\mathbb{R}/\mathbb{G}}T_{c^{\prime}}1_{\mu,c},-)$. We also have $\mathfrak{A}_{c}\circ\mathfrak{B}_{c}(M)\cong_{\mathrm{a}}M$ for any $M$.
@@
$\mu^{\mathbb{G}}_{<c}(*)\ni 1_{\mu,c}:=\bigoplus_{c^{\prime}\in\mathbb{G}}\mathbb{K}_{\left\{(s,t)\mathrel{}\middle|\mathrel{}0<s<c,c^{\prime}\leq t<s+c^{\prime}\right\}}.$ (6.14)
By mimicking the argument in [Kuw24] and Lemma 5.2, one can see that the endmorphism is almost isomorphic to $L_{0}^{\mathbb{G}}/T^{c}L_{0}^{\mathbb{G}}$. We then have a functor
$\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{c}L_{0}^{\mathbb{G}})\rightarrow\mu^{\mathbb{G}}_{<c}(*);M\mapsto M\otimes 1_{\mu,c}.$ (6.15)
Then this is almost fully faithful. The claim about the adjoint can be proved in the same way as Proposition 5.7. ∎
@@
One can see that this is right orthogonal to $1_{\mu,1}$. Indeed, if one apply $\operatorname{Hom}(1_{\mu,1},-)$ to the sequence the resulting sequence is
$\cdots\xrightarrow{T^{1/2}}\Lambda_{0}/T\Lambda_{0}\xrightarrow{T^{1/2}}\Lambda_{0}/T\Lambda_{0}\xrightarrow{T^{1/2}}\Lambda_{0}/T\Lambda_{0}\xrightarrow{T^{1/2}}\cdots,$ (6.17)
which is acyclic. Hence $1_{\mu,1}$ does not generate the whole category. The author would like to thank T. Asano for asking a related question on the draft version.
@@
$\mu^{\mathbb{G}}_{<c}(T^{*}N)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(N,\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{c}L_{0}^{\mathbb{G}})).$ (6.18)
@@
$\mu^{\mathbb{R}}_{<c}(T^{*}X)\hookleftarrow_{\mathrm{a}}\mathrm{Sh}(X,\Lambda_{0}/T^{c}\Lambda_{0}).$ (6.19)
Proof.
For any $\mathcal{E},\mathcal{F}\in\mathrm{Sh}(N,\mathrm{Mod}(L_{0}^{\mathbb{G}}/T^{c}L_{0}^{\mathbb{G}}))$, we have
$\operatorname{Hom}(\mathfrak{B}_{c}(\mathcal{E}),\mathfrak{B}_{c}(\mathcal{F}))\cong\operatorname{Hom}(\mathcal{E},\mathfrak{A}_{c}\circ\mathfrak{B}_{c}(\mathcal{F})).$ (6.20)
The RHS is almost isomorphic to $\operatorname{Hom}(\mathcal{E},\mathcal{F})$, since $\mathfrak{A}_{c}\circ\mathfrak{B}_{c}(\mathcal{F})\cong_{\mathrm{a}}\mathcal{F}$. ∎
6.5 Variant 2: Higher-dimensional version
@@
For a proper convex cone $\gamma$ in $\mathbb{R}^{n}$, we say that $\gamma$ is simplicial if there exists a linear isomorphism of $\mathbb{R}^{n}$ which gives an isomorphism $\gamma\cong\mathbb{R}^{n}_{\geq 0}$.
Let $\gamma$ be a simplicial closed polyhedral proper convex cone in $\mathbb{R}^{n}$ with nonempty interior. Then $\gamma$ has a semigroup structure with respect to the addition. We denote the corresponding polynomial ring by $\mathbb{K}[\gamma]$. We denote the indeterminate corresponding to $a\in\gamma$ by $T^{a}$. Let $|\cdot|$ be the Euclidean norm of $\mathbb{R}^{n}$. For $r\in\mathbb{R}_{>0}$, we denote the ideal of $\mathbb{K}[\gamma]$ generated by $T^{a}$’s with $|a|>r$ by $\mathfrak{m}(r)$. We set
@@
[…]
By tensoring Theorem 6.3 $n$ times, we get the following:
@@
$\mu^{\gamma}(T^{*}X)\hookrightarrow_{\mathrm{a}}\mathrm{Sh}(X,\Lambda_{0}^{\gamma}).$ (6.23)
@@ 7.1 Non-conic microsupport
Let $U$ be an open subset of $X$. Let $\phi$ be a continuous function on $U$ which is bounded below. For any connected open subset $V\subset U$, we denote the infimum value of $\phi$ by $\phi_{V}$. We define a sheaf $R^{\phi}$ as follows: For any connected open subset $V$, we set $R^{\phi}(V):=R$. For an open inclusion of connected open subsets $W\subset V$, we have $\phi_{W}\geq\phi_{V}$, we set a structure morphism $R^{\phi}(V)\rightarrow R^{\phi}(W)$ by $T^{\phi_{W}-\phi_{V}}$. Sheafifying this, we get a sheaf on $U$. We also set $R^{\phi}_{c}:=R^{\phi}\otimes_{R}R_{c}$ for $c\geq 0$.
@@
For $\mathcal{E}\in\mathrm{Sh}(X,R)$, we set $\mu\mathrm{supp}(\mathcal{E})$ to be the closure of the complement of the following set:
[…]
@@ 7.2 Relation to usual microsupport
Since an object of $\mathrm{Sh}(X,R)$ is a sheaf on a manifold, we can also define the usual microsupport of [KS94].
@@
For a subset $A\subset T^{*}X$, we set
$\mathbb{R}_{\geq 0}\cdot A:=\left\{(x,\xi)\in T^{*}X\mathrel{}\middle|\mathrel{}(x,\xi^{\prime})\in A,c\in\mathbb{R}_{\geq 0}\text{ s.t.}(x,\xi)=(x,c\xi^{\prime})\right\}$ (7.4)
@@
[…]